OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6

OpenStudy (campbell_st):

distribute the 3 and -2 3x + 3 = -2x +2 + 6 simplify 3x + 3 = -2x + 8 add 2x to both sides 5x + 3 = 8 subtract 3 from both sides 5x = 5 now you can solve for x.

OpenStudy (anonymous):

thank you